OpenStudy (anonymous):

Estimate the following using clustering: 21.8 + 20.1 + 19.5 + 19.2 + 19.1 (Points : 1) 99.7 100 95 20

OpenStudy (ybarrap):

All numbers are clustering around 20.0 So how can you use that?

OpenStudy (ybarrap):

If each term can be approximated by 20 and you have 5 terms, what is the sum of this approximation?

OpenStudy (anonymous):

i dont know sorry

OpenStudy (anonymous):

so the answer will be 20

OpenStudy (amoodarya):

21.8 + 20.1 + 19.5 + 19.2 + 19.1= (20+1.8)+(20+.1)+(20-.5)+(20-.8)+(20-.9)= 5*20 +(1.8+0.1-0.5-0.8-0.9)

OpenStudy (anonymous):

99.7